1. Identify and describe the five most important ratios for a potential bond investor in a corporate bond issuance. Explain the reasoning for your selection. Illustrate your discussion using numerical examples.
Answer #1

1) Leverage Ratio:-

It is important to understand the risk profile of the organization before investing in their bond. To get an idea about that, we calculate Leverage Ratio.

Too much debt or too little debt both are bad for an organization. Too little debt can make an opportunity loss for shareholders to earn more. For other cases, it can be risky for the organization to sustain for a longer period.  

There are some leverage ratios to calculate,

Debt\: Ratio = \frac{Debt}{Total\, capital\, Employed}

Debt\: Equity\: Ratio = \frac{Total Debt}{Net\: Worth}

2) Current Ratio:-

It gives an idea of the company's liquidity. This is extremely important for a company to be able to meet its obligations on time. This ratio will show you that. You can compare that company's current ratio to the industry average to get an idea.

Current\: Ratio = \frac{Current\, Assets}{Current \:Liabilities}

3) Assets Turnover:-

Assets are bought to generate incomes for the organization. This ratio shows how efficiently the organization using its assets. Because if assets don't generate incomes then it can block huge money in the business.

Asset\: Turnover\: Ratio = \frac{Sales}{Net \:Assets}

4) Profit Margin:-

A company needs to earn profit to sustain. As an investor one should access a company's profitability ratios.

Net\: Profit\: Margin = \frac{Net Profit}{Sales}

Operating\: Profit\: Margin = \frac{EBIT}{Sales}

EBIT = Earning Before Interest Tax

Net Margin Based on NOPAT is also a very important ratio to calculate. PAT excludes interest on borrowing. NOPAT take care all of these.

Net\: Profit\: Margin(using\, NOPAT)= \frac{EBIT(1-T)}{Sales}=\frac{NOPAT}{Sales}

5) Price Earning Ratio(P/E Ratio):-

It is widely used by security analysts to value a firm's performance as expected by investors. This is like a market appraisal for the firms.

P/E \: Ratio= \frac{Market\, Value\, per\, Share}{Earning Per Share(EPS)}

These are the ratios one need to take care when accessing a firm and buying a bond.